# Consistent eccentricities for gravitational wave astronomy: Resolving discrepancies between astrophysical simulations and waveform models

### v2

@@ III.4 Implications for existing work
A number of works have ascertained the detectability of eccentric mergers within the LVK band as well as in other detector bands, using data from astrophysical simulations such as CMC. As we have shown in the previous sections, the waveform eccentricities are not in agreement with the eccentricities given by simulations. Hence, we would expect these detectability estimates to change when defining eccentricity consistently in the simulations. For instance, Ref. Zevin et al. (2021b) found that $\approx$60of potentially detectable eccentric sources with $e\gtrsim 0.2$ would be missed due to non-inclusion of eccentric templates in search pipelines. Using the same dataset as Ref. Zevin et al. (2021b) and changing the eccentricities from the $e_{\rm W03}$ definition to the $e_{t}^{\rm 2PN}$ definition, we find that only 30% of potentially detectable eccentric sources with $e\gtrsim 0.2$ are missed—a discrepancy by a factor of $\approx 2$. The decrease in the number of missed signals is due to the shifting of the histogram of eccentricities towards lower values, which are better approximated by non-eccentric templates and hence less likely to be selected against. As another illustration, we calculate limits on the cluster branching fraction $\beta_{c}$ i.e the fraction of events in the total population coming from clusters using the same method and dataset as used in Ref. Zevin et al. (2021a). We find that assuming $N_{\rm ecc}=0$ eccentric detections in GWTC-3, we would have inferred $\beta_{c}<86\%$ ($90\%$ credible level) using the $e_{\rm W03}$ estimates, and infer $\beta_{c}<63\%$ using the $e_{t}^{\rm 2PN}$ estimates. On the other hand, assuming $N_{\rm ecc}=1$ gives $\beta_{c}>10\%$ using the $e_{\rm W03}$ estimates and $\beta_{c}>8\%$ using the $e_{t}^{\rm 2PN}$ estimates.
@@
Similarly, any study that uses the detected eccentricities (or lack thereof) to place constraints on the overall merger rate in eccentric systems in astrophysical environments will also be affected by these discrepancies. For instance, Ref. Abac et al. (2023) use the non-detection of orbital eccentricity from a waveform-independent search to place an upper limit of $0.34\,{\rm Gpc^{-3}\,{yr}^{-1}}$ on the merger rate of binaries in dense stellar clusters, assuming a population model from CMC simulations. This population model is taken from Ref. Zevin et al. (2021b) with eccentricities defined using the W03 prescription, with additional conditioning on the mass ($M>70$) and eccentricity ($e<0.3$) parameter ranges of their search. From Figure 2, we see that the relative number of mergers having $e<0.3$ increases when accounting for the prescription in this work. This means that an upper limit derived using the $e_{t}^{\rm 2PN}$ estimates from this work would be lower than the ones quoted in Ref. Abac et al. (2023).
@@
We also note that Ref. Romero-Shaw et al. (2022), which uses signatures of non-zero eccentricity in four GW events to constrain the fraction of BBH mergers formed dynamically, does account for the disparate conventions for eccentricity while making their comparisons. They do so in an approximate fashion, similar to the calculation of $e_{\rm W03}^{f_{22}}$ described in Section III.1. We do not expect conclusions reported there to change significantly.
IV Conclusion
@@
Orbital eccentricity in a compact binary is believed to be a telltale sign of its formation history. Consequently, there has been progress on building waveform models for eccentric sources as well as on simulating expectations from astrophysical formation channels. In this work, we note that the conventions used to define eccentricity in the GW waveform and astrophysical simulation communities differ from each other. The major difference in these conventions lies in the choice of the reference frequency at which eccentricity is defined—GW waveforms typically define eccentricity at a fixed orbit averaged 22-mode GW frequency, while astrophysical simulations define eccentricity at a fixed peak GW frequency following the W03 prescription (described in Section II.2). Further, we delineate an eccentricity calculation procedure that can be applied to data from astrophysical simulations to extract eccentricities consistent with conventions laid down by the GW waveform community. This procedure involves evolution of ordinary differential equations describing the time-evolution of eccentricity and frequency obtained from PN expansions, and extracting eccentricity consistently at the orbit-averaged 22 mode frequency.
@@
Using compact binaries from the publicly-available CMC catalog as an illustration, we find that there is a discrepancy between eccentricity estimates derived using the W03 prescription and the one described in this work. The discrepancy is small for $e\lesssim 0.1$, but is large for higher values of eccentricity. Notably, all binaries with $e_{\rm W03}\approx 1$ (thought to be GW captures happening in the LVK band) have $e_{t}^{\rm 2PN}$ between 0.40.40.40.40.80.80.80.8. We investigate the effect that adding 2PN corrections to the leading order evolution has on derived eccentricities, and also the effects of not including spins in our prescription. We also compare our $e_{t}^{\rm 2PN}$ estimates to $e_{\rm gw}$ estimates derived from the gw_eccentricity package using the SEOBNRv4EHM waveform. We find that the $e_{t}^{\rm 2PN}$ estimates should be sufficiently accurate in extracting eccentricities, and also computationally efficient to implement as a post-processing step in simulations. However, as outlined in Ref. Shaikh et al. (2023), gw_eccentricity is the most accurate way of extracting eccentricity and should be used when high-precision eccentricity estimates are necessary.
@@
The different conventions for defining eccentricity currently employed in the literature mean that one should be careful in comparing eccentricity estimates/upper limits derived from real data to astrophysical simulations, creating mock populations of eccentric sources based on astrophysical expectations, and other investigations that pair predictions from astrophysical models with GW data analysis. As an illustration, we show that the detectability estimates of eccentric sources and branching fractions from Ref. Zevin et al. (2021b) will change when ensuring a consistent definition of eccentricity. We recommend that catalogs created from astrophysical simulations include $e_{t}^{\rm 2PN}$ estimates (extracted at a constant $Mf_{22}$) for clarity and ease of direct comparison to GW observations.
@@ Acknowledgements
We are grateful to Isobel Romero-Shaw for a thorough reading of our draft and useful suggestions. We thank Kaushik Paul for insighful inputs regarding PN equations and sharing code for the same, and Antoni Ramos-Buades for providing access to the SEOBNRv4EHM approximant. We also thank Vijay Varma, Prayush Kumar, Bala Iyer, and Akash Maurya for useful discussions. Example python scripts are available at https://github.com/adivijaykumar/standard-eccentricity-scripts, and data products are available on Zenodo Vijaykumar et al. (2024).
AV is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference number 568580), and the Department of Atomic Energy, Government of India, under Project No. RTI4001. AV also acknowledges support by a Fulbright Program grant under the Fulbright-Nehru Doctoral Research Fellowship, sponsored by the Bureau of Educational and Cultural Affairs of the United States Department of State and administered by the Institute of International Education and the United States-India Educational Foundation. AGH is supported by NSF grants AST-2006645 and PHY2110507. AGH gratefully acknowledges the ARCS Foundation Scholar Award through the ARCS Foundation, Illinois Chapter with support from the Brinson Foundation. MZ gratefully acknowledges funding from the Brinson Foundation in support of astrophysics research at the Adler Planetarium. This work has made use of the numpy Harris et al. (2020), scipy Virtanen et al. (2020), matplotlib Hunter (2007), astropy Astropy Collaboration et al. (2013, 2018), jupyter Kluyver et al. (2016), pandas McKinney (2010), seaborn Waskom (2021), and lalsuite LIGO Scientific Collaboration et al. (2018) software packages.
@@ Appendix A PN formulae for eccentricity and velocity evolution
$\alpha_{2}$ $=\frac{125361e^{4}+718008e^{2}-28\left(3328e^{4}+23259e^{2}+8168\right)\eta-67608}{168(121e+304)\epsilon^{2}}$ (17)
$\alpha_{3}$ $=\frac{1970\pi v^{3}\epsilon^{5}}{121e+304}\phi_{e}^{\rm rad}$ (18)
@@
$\mathrel{\phantom{=}}-16\left(-284256\eta^{2}-801495\eta+168336(2\eta-5)\epsilon+952397\right)]$ (19)
@@
$\beta_{2}$ $=\dfrac{e^{6}(11717-8288\eta)+e^{4}(171038-141708\eta)-120e^{2}(1330\eta-731)+16(743-924\eta)}{56\left(37e^{4}+292e^{2}+96\right)\epsilon^{2}}$ (20)
$\beta_{3}$ $=\frac{768\pi v^{3}\epsilon^{7}}{74e^{4}+584e^{2}+192}\phi_{e}$ (21)
@@
$\mathrel{\phantom{=}}+32\left(-59472\eta^{2}-141093\eta+9072(2\eta-5)\epsilon+11257\right)]\mbox{\quad.\quad}$ (22)
Here, $\epsilon=\sqrt{1-e^{2}}$. Here, the “enhancement factors” $\phi_{e}$, $\tilde{\phi}_{e}$, and $\phi_{e}^{\rm rad}$ are given by
$\phi_{e}$ $=\frac{\frac{428340e^{12}}{9958749469}-\frac{5034498e^{10}}{7491716851}+\frac{9293260e^{8}}{3542508891}+\frac{48176523e^{6}}{177473701}+\frac{157473274e^{4}}{30734301}+\frac{18970894028e^{2}}{2649026657}+1}{\epsilon^{10}}$ (23)
$\tilde{\phi}_{e}$ $=\frac{-\frac{328675e^{10}}{8674876481}-\frac{4679700e^{8}}{6316712563}-\frac{2640201e^{6}}{993226448}+\frac{37570495e^{4}}{98143337}+\frac{413137256e^{2}}{136292703}+1}{\epsilon^{9}}$ (24)
$\phi_{e}^{\rm rad}$ $=\frac{192\epsilon(\epsilon\phi_{e}-\tilde{\phi}_{e})}{985e^{2}}\mbox{\quad.\quad}$ (25)
@@ Appendix B Variation in gw_eccentricity estimates calculated from different waveform approximants
The prescription of calculating eccentricity $e_{\rm gw}$, proposed by Shaikh et al. (2023) and implemented in the gw_eccentricity python package only uses the GW waveform as seen by the detector as input, as described in Section III.3. This definition is beneficial since eccentricity is now linked to a direct observable. However, different waveforms will, in general, yield different values of $e_{\rm gw}$ for the same binary. This could be due to a variety of reasons: differences in the modeling of the dynamics, different internal definitions of eccentricity in the model, or even choices made while implementing the approximant in some codebase such as LALSimulation LIGO Scientific Collaboration et al. (2018). Figure 6 illustrates this point by comparing $e_{\rm gw}$ derived from two waveforms, EccentricTD Tanay et al. (2016) and SEOBNRv4EHM Ramos-Buades et al. (2022). For this illustration, we assume an equal mass binary with a total mass of $M=50\,M_{\odot}$. We vary the initial eccentricity $e_{0}$ (defined at initial frequency $f_{0}=5\,{\rm Hz}$) supplied to the model between $10^{-2}$ and $0.8$, and extract $e_{\rm gw}$ from both EccentricTD $\quantity(e_{\rm gw}^{\rm EccTD})$ and SEOBNRv4EHM $\quantity(e_{\rm gw}^{\rm SEOB})$, at both the initial frequency $f_{0}$ and also at $f_{\rm ref}=20\,{\rm Hz}$. We see that $e_{\rm gw}^{\rm SEOB}(f_{0})$ matches all input values $e_{0}$, whereas $e_{\rm gw}^{\rm EccTD}(f_{0})$ differs from $e_{0}$ by a few percent. This is in agreement with investigations performed in Ref. Shaikh et al. (2023), where they found that the $e_{\rm gw}^{\rm SEOB}\quantity(f_{0})=e_{0}$ trend is maintained for all $e_{0}>10^{-5}$, whereas $e_{\rm gw}^{\rm EccTD}$ deviates from the $e_{\rm gw}^{\rm EccTD}\quantity(f_{0})=e_{0}$ trend. We also see that there is a large discrepancy (factor of a few) between $e_{\rm gw}^{\rm EccTD}(f_{\rm ref})$ and $e_{\rm gw}^{\rm SEOB}(f_{\rm ref})$, strengthening the evidence for differences in the ways each of these waveform approximants define eccentricity. We use the SEOBNRv4EHM approximant for our comparisons in Section III.3.

### v3

@@ 3.4 Implications for existing work
A number of works have ascertained the detectability of eccentric mergers within the LVK band as well as in other detector bands, using data from astrophysical simulations such as CMC. As we have shown in the previous sections, the waveform eccentricities are not in agreement with the eccentricities given by simulations. Hence, we would expect these detectability estimates to change when defining eccentricity consistently in the simulations. For instance, Zevin et al. (2021a) found that $\approx$60of potentially detectable eccentric sources with $e\gtrsim 0.2$ would be missed due to non-inclusion of eccentric templates in search pipelines. Using the same dataset as Zevin et al. (2021a) and changing the eccentricities from the $e_{\rm W03}$ definition to the $e_{t}^{\rm 2PN}$ definition, we find that only 30% of potentially detectable eccentric sources with $e\gtrsim 0.2$ are missed—a discrepancy by a factor of $\approx 2$. The decrease in the number of missed signals is due to the shifting of the histogram of eccentricities towards lower values, which are better approximated by non-eccentric templates and hence less likely to be selected against. As another illustration, we calculate limits on the cluster branching fraction $\beta_{c}$ i.e the fraction of events in the total population coming from clusters using the same method and dataset as used in Zevin et al. (2021b). We find that assuming $N_{\rm ecc}=0$ eccentric detections in GWTC-3, we would have inferred $\beta_{c}<86\%$ ($90\%$ credible level) using the $e_{\rm W03}$ estimates, and infer $\beta_{c}<63\%$ using the $e_{t}^{\rm 2PN}$ estimates. On the other hand, assuming $N_{\rm ecc}=1$ gives $\beta_{c}>10\%$ using the $e_{\rm W03}$ estimates and $\beta_{c}>8\%$ using the $e_{t}^{\rm 2PN}$ estimates.
@@
Similarly, any study that uses the detected eccentricities (or lack thereof) to place constraints on the overall merger rate in eccentric systems in astrophysical environments will also be affected by these discrepancies. For instance, Abac et al. (2023) use the non-detection of orbital eccentricity from a waveform-independent search to place an upper limit of $0.34\,{\rm Gpc^{-3}\,{yr}^{-1}}$ on the merger rate of binaries in dense stellar clusters, assuming a population model from CMC simulations. This population model is taken from Zevin et al. (2021a) with eccentricities defined using the W03 prescription, with additional conditioning on the mass ($M>70$) and eccentricity ($e<0.3$) parameter ranges of their search. From Figure 2, we see that the relative number of mergers having $e<0.3$ increases when accounting for the prescription in this work. This means that an upper limit derived using the $e_{t}^{\rm 2PN}$ estimates from this work would be lower than the ones quoted in Abac et al. (2023).
@@
We also note that Romero-Shaw et al. (2022), which uses signatures of non-zero eccentricity in four GW events to constrain the fraction of BBH mergers formed dynamically, does account for the disparate conventions for eccentricity while making their comparisons. They do so in an approximate fashion, similar to the calculation of $e_{\rm W03}^{f_{22}}$ described in Section 3.1. We do not expect conclusions reported there to change significantly.
4 Conclusion
@@
Orbital eccentricity in a compact binary is believed to be a telltale sign of its formation history. Consequently, there has been progress on building waveform models for eccentric sources as well as on simulating expectations from astrophysical formation channels. In this work, we note that the conventions used to define eccentricity in the GW waveform and astrophysical simulation communities differ from each other. The major difference in these conventions lies in the choice of the reference frequency at which eccentricity is defined—GW waveforms typically define eccentricity at a fixed orbit averaged 22-mode GW frequency, while astrophysical simulations define eccentricity at a fixed peak GW frequency following the W03 prescription (described in Section 2.2). Further, we delineate an eccentricity calculation procedure that can be applied to data from astrophysical simulations to extract eccentricities consistent with conventions laid down by the GW waveform community. This procedure involves evolution of ordinary differential equations describing the time-evolution of eccentricity and frequency obtained from PN expansions, and extracting eccentricity consistently at the orbit-averaged 22 mode frequency.
@@
Using compact binaries from the publicly-available CMC catalog as an illustration, we find that there is a discrepancy between eccentricity estimates derived using the W03 prescription and the one described in this work. The discrepancy is small for $e\lesssim 0.1$, but is large for higher values of eccentricity. Notably, all binaries with $e_{\rm W03}\approx 1$ (thought to be GW captures happening in the LVK band) have $e_{t}^{\rm 2PN}$ between 0.40.40.40.40.80.80.80.8. We investigate the effect that adding 2PN corrections to the leading order evolution has on derived eccentricities, and also the effects of not including spins in our prescription. We also compare our $e_{t}^{\rm 2PN}$ estimates to $e_{\rm gw}$ estimates derived from the gw_eccentricity package using the SEOBNRv4EHM waveform. We find that the $e_{t}^{\rm 2PN}$ estimates should be sufficiently accurate in extracting eccentricities, and also computationally efficient to implement as a post-processing step in simulations. However, as outlined in Shaikh et al. (2023), gw_eccentricity is the most accurate way of extracting eccentricity and should be used when high-precision eccentricity estimates are necessary.
@@
The different conventions for defining eccentricity currently employed in the literature mean that one should be careful in comparing eccentricity estimates/upper limits derived from real data to astrophysical simulations, creating mock populations of eccentric sources based on astrophysical expectations, and other investigations that pair predictions from astrophysical models with GW data analysis. As an illustration, we show that the detectability estimates of eccentric sources and branching fractions from Zevin et al. (2021a) will change when ensuring a consistent definition of eccentricity. We recommend that catalogs created from astrophysical simulations include $e_{t}^{\rm 2PN}$ estimates (extracted at a constant $Mf_{22}$) for clarity and ease of direct comparison to GW observations.
@@ Acknowledgements
We are grateful to Isobel Romero-Shaw for a thorough reading of our draft and useful suggestions. We thank Kaushik Paul for insighful inputs regarding PN equations and sharing code for the same, and Antoni Ramos-Buades for providing access to the SEOBNRv4EHM approximant. We also thank Vijay Varma, Prayush Kumar, Bala Iyer, and Akash Maurya for useful discussions. Example python scripts (Vijaykumar et al., 2024a) and data products (Vijaykumar et al., 2024b) are available on Zenodo.
AV is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference number 568580), and the Department of Atomic Energy, Government of India, under Project No. RTI4001. AV also acknowledges support by a Fulbright Program grant under the Fulbright-Nehru Doctoral Research Fellowship, sponsored by the Bureau of Educational and Cultural Affairs of the United States Department of State and administered by the Institute of International Education and the United States-India Educational Foundation. AGH is supported by NSF grants AST-2006645 and PHY2110507. AGH gratefully acknowledges the ARCS Foundation Scholar Award through the ARCS Foundation, Illinois Chapter with support from the Brinson Foundation. MZ gratefully acknowledges funding from the Brinson Foundation in support of astrophysics research at the Adler Planetarium.
[…]
@@ Appendix A PN formulae for eccentricity and velocity evolution
$\alpha_{2}$ $=\frac{125361e^{4}+718008e^{2}-28\left(3328e^{4}+23259e^{2}+8168\right)\eta-67608}{168(121e+304)\epsilon^{2}}$ (A4)
$\alpha_{3}$ $=\frac{1970\pi v^{3}\epsilon^{5}}{121e+304}\phi_{e}^{\rm rad}$ (A5)
@@
$\mathrel{\phantom{=}}-16\left(-284256\eta^{2}-801495\eta+168336(2\eta-5)\epsilon+952397\right)]$ (A6)
@@
$\beta_{2}$ $=\dfrac{e^{6}(11717-8288\eta)+e^{4}(171038-141708\eta)-120e^{2}(1330\eta-731)+16(743-924\eta)}{56\left(37e^{4}+292e^{2}+96\right)\epsilon^{2}}$ (A7)
$\beta_{3}$ $=\frac{768\pi v^{3}\epsilon^{7}}{74e^{4}+584e^{2}+192}\phi_{e}$ (A8)
@@
$\mathrel{\phantom{=}}+32\left(-59472\eta^{2}-141093\eta+9072(2\eta-5)\epsilon+11257\right)]\mbox{\quad.\quad}$ (A9)
Here, $\epsilon=\sqrt{1-e^{2}}$ and the “enhancement factors” $\phi_{e}$, $\tilde{\phi}_{e}$, and $\phi_{e}^{\rm rad}$ are given by
$\phi_{e}$ $=\frac{\frac{428340e^{12}}{9958749469}-\frac{5034498e^{10}}{7491716851}+\frac{9293260e^{8}}{3542508891}+\frac{48176523e^{6}}{177473701}+\frac{157473274e^{4}}{30734301}+\frac{18970894028e^{2}}{2649026657}+1}{\epsilon^{10}}$ (A10)
$\tilde{\phi}_{e}$ $=\frac{-\frac{328675e^{10}}{8674876481}-\frac{4679700e^{8}}{6316712563}-\frac{2640201e^{6}}{993226448}+\frac{37570495e^{4}}{98143337}+\frac{413137256e^{2}}{136292703}+1}{\epsilon^{9}}$ (A11)
$\phi_{e}^{\rm rad}$ $=\frac{192\epsilon(\epsilon\phi_{e}-\tilde{\phi}_{e})}{985e^{2}}\mbox{\quad.\quad}$ (A12)
@@ Appendix B Variation in gw_eccentricity estimates calculated from different waveform approximants
The prescription of calculating eccentricity $e_{\rm gw}$, proposed by Shaikh et al. (2023) and implemented in the gw_eccentricity python package only uses the GW waveform as seen by the detector as input, as described in Section 3.3. This definition is beneficial since eccentricity is now linked to a direct observable. However, different waveforms will, in general, yield different values of $e_{\rm gw}$ for the same binary. This could be due to a variety of reasons: differences in the modeling of the dynamics, different internal definitions of eccentricity in the model, or even choices made while implementing the approximant in some codebase such as LALSimulation (LIGO Scientific Collaboration et al., 2018). Figure 6 illustrates this point by comparing $e_{\rm gw}$ derived from two waveforms, EccentricTD (Tanay et al., 2016) and SEOBNRv4EHM (Ramos-Buades et al., 2022). For this illustration, we assume an equal mass binary with a total mass of $M=50\,M_{\odot}$. We vary the initial eccentricity $e_{0}$ (defined at initial frequency $f_{0}=5\,{\rm Hz}$) supplied to the model between $10^{-2}$ and $0.8$, and extract $e_{\rm gw}$ from both EccentricTD $\quantity(e_{\rm gw}^{\rm EccTD})$ and SEOBNRv4EHM $\quantity(e_{\rm gw}^{\rm SEOB})$, at both the initial frequency $f_{0}$ and also at $f_{\rm ref}=20\,{\rm Hz}$. We see that $e_{\rm gw}^{\rm SEOB}(f_{0})$ matches all input values $e_{0}$, whereas $e_{\rm gw}^{\rm EccTD}(f_{0})$ differs from $e_{0}$ by a few percent. This is in agreement with investigations performed in Shaikh et al. (2023), where they found that the $e_{\rm gw}^{\rm SEOB}\quantity(f_{0})=e_{0}$ trend is maintained for all $e_{0}>10^{-5}$, whereas $e_{\rm gw}^{\rm EccTD}$ deviates from the $e_{\rm gw}^{\rm EccTD}\quantity(f_{0})=e_{0}$ trend. We also see that there is a large discrepancy (factor of a few) between $e_{\rm gw}^{\rm EccTD}(f_{\rm ref})$ and $e_{\rm gw}^{\rm SEOB}(f_{\rm ref})$, strengthening the evidence for differences in the ways each of these waveform approximants define eccentricity. We use the SEOBNRv4EHM approximant for our comparisons in Section 3.3.